OpenStudy (anonymous):

Anyone up for this one - The average height of a boy in the U.S., from birth thru 60 months, can be modeled by y=2.9√x+20.1 where y is the average height in inches of boys who are x months of age. What would the expected difference in height be between a child 49 months and a child 16 months?

OpenStudy (dumbcow):

if x=49, y = (2.9)(7)+20.1 = 40.4 if x=16, y = (2.9)(4) + 20.1 = 31.7 difference = 40.4 - 31.7

OpenStudy (anonymous):

so 8.7 thank you